# Schema Generation for Large Knowledge Graphs Using Large Language Models

### v1

@@ Abstract
Schemas are vital for ensuring data quality in the Semantic Web and natural language processing. Traditionally, their creation demands substantial involvement from knowledge engineers and domain experts. Leveraging the impressive capabilities of large language models (LLMs) in related tasks like ontology engineering, we explore automatic schema generation using LLMs. To bridge the resource gap, we introduce two datasets: YAGO Schema and Wikidata EntitySchema, along with evaluation metrics. The LLM-based pipelines effectively utilize local and global information from knowledge graphs (KGs) to generate validating schemas in Shape Expressions (ShEx). Experiments demonstrate LLMs’ strong potential in producing high-quality ShEx schemas, paving the way for scalable, automated schema generation for large KGs. Furthermore, our benchmark introduces a new challenge for structured generation, pushing the limits of LLMs on syntactically rich formalisms.
@@ 1 Introduction
Validating KGs against predefined schemas is a crucial technique for quality assessment, effective in identifying structural and semantic inconsistencies to maintain data integrity (Gayo et al., 2018; Scherp et al., 2024). These validating schemas define the expected rules and patterns for the data, focusing on specific sets of nodes (Ahmetaj et al., 2025). To represent these schemas, distinct languages including W3C standards like Shapes Constraint Language (SHACL) (Knublauch and Kontokostas, 2017), Shape Expressions (ShEx) (Prud’hommeaux et al., 2014), and PG-Schema (Angles et al., 2023) have emerged. Developing high-quality validation schemas aligned with user needs is a considerable challenge (Rabbani et al., 2022). Current automatic schema generation via pattern aggregation often inherits noise and errors from the source KGs (Rabbani et al., 2022). As a result, the primary approach involves manually writing schemas or refining those produced by basic pattern mining—a process that is both time-consuming and expensive, particularly for KGs with potentially millions of classes (Rabbani et al., 2022). The demonstrated success of LLMs in areas like ontology engineering (Lo et al., 2024; Zhang et al., 2025) and graph-based reasoning (Wang et al., 2023) highlights their proficiency with structured data. Specifically, LLMs show aptitude for processing structured information, reasoning across graph and text modalities, and performing structured generation (He et al., 2024b; Jin et al., 2024). This aptitude strongly suggests their potential for automatically generating KG validation schemas.
Building on this, this paper investigates how LLMs can undertake this automatic schema generation for large KGs, especially where such schemas are missing or incomplete. Specifically, we address the research question: How can LLMs generate high-quality ShEx shapes for a given KG? We identify key challenges inherent to this task. First, due to the sheer scale of large KGs, pinpointing the essential information required for accurate ShEx generation by LLMs is difficult. Second, even when relevant information is identified, investigating how LLMs can effectively leverage this structured data for schema generation is required. Finally, the relative novelty and ongoing development of the ShEx standard mean that publicly available ShEx corpora are limited, potentially hindering LLM familiarity with the required syntax and conventions. To address these challenges, this paper makes the following main contributions:
@@
We introduce a benchmark comprising two datasets derived from large KGs—YAGO and Wikidata—along with associated metrics designed for comprehensive evaluation from diverse perspectives.
We propose and evaluate pipelines designed to efficiently generate high-quality ShEx schemas by leveraging the structured generation capabilities of LLMs111We will release our code upon acceptance..
@@ 2 Related Work
Schema Generation
KG schemas are broadly classified into semantic, validating, and emergent schemas (Hogan et al., 2021). We focus on validating schemas, typically defined using shapes. A shape identifies a set of focus nodes within a data graph and specifies constraints they must adhere to. ShEx is a prominent language for expressing such schemas, especially in Wikidata.
@@
Structured Generation
Structured generation, or constrained decoding, enables LLMs to produce outputs adhering to precise formats, vital for applications like code generation (Ugare et al., 2024) and tool calling (Zhang et al., 2024). This involves generating token sequences that satisfy specified constraints, often defined by regular expressions (regex) or context-free grammars (CFGs). A key challenge is efficiently applying these constraints over the LLM’s large vocabulary without degrading speed or performance (Dong et al., 2024; Koo et al., 2024). Fine-tune-free methods primarily fall into two categories: token-level constrained decoding and prompt-level output structuring. Token-level methods directly influence token selection during decoding. They often compile regex and CFGs into automata (finite state machines or pushdown automata) that operate in the token space (Koo et al., 2024). This generation efficiency is further enhanced through techniques like states tracking and indexing (Willard and Louf, 2023), and token categorization and caching (Dong et al., 2024). Prompt-level methods leverage LLMs’ instruction-following or function-calling capabilities to guide the overall output structure, which is subsequently parsed and validated. Instructor (Liu and Contributors, 2024) exemplifies this by converting Pydantic models into LLM-understandable schemas (e.g., JSON Schema) and incorporating a retry mechanism with feedback.
@@ 3 Problem Formulation
Schema generation is a challenging task for both traditional rule-based methods (Rabbani et al., 2023) and ML approaches like LLMs, primarily due to several factors. Large KGs often contain quality issues, and pattern extraction may inadvertently encode these issues into generated shapes. Moreover, there is a lack of benchmarks and high-quality ShEx ground truths, making training and evaluation for ML-based approaches particularly difficult.
@@ 4.1 Dataset
To address the need for benchmarks, we introduce two new dataset, YAGO Schema (YAGOS) and Wikidata EntitySchema (WES). Together, comprising 86 ShEx schema with a total of 2,552 constraints. Detailed statistics are presented in Table 1, and the dataset construction process is outlined in Appendix A. Each schema in our datasets targets a specific class within the respective KG. While ShEx offers a rich syntax, the schemas in our datasets employ a community-prevalent subset to enhance simplicity and readability while remaining functional (detailed specification in Appendix B). In our setting, to simplify the syntax representation and ease evaluation and comparison, a shape schema includes a start shape that targets the focus class. The remaining shapes in the schema serve as references, each describing one or more classes that specify the range of objects allowed for certain predicates in the start shape. The schemas predominantly use four fundamental node constraint types: node kind constraints (e.g., IRI), datatype constraints (e.g., xsd:decimal, rdf:langString), value set constraints (often specifying object values, e.g., [schema:Organization]), and shape references (e.g., @<Person>, which links to another defined shape like <Person> { rdf:type [ schema:Person ] }).
@@ 4.2.1 Similarity Metrics
Given that automatically generated schemas often require manual refinement (Rabbani et al., 2022), quantifying the structural similarity between a generated schema and its ground truth counterpart is crucial. To facilitate this comparison, we model each shape schema as a rooted, labeled tree graph. The shape label (or focus node) serves as the root, predicates form the first level of child nodes, node constraints linked to predicates occupy the next level, and cardinalities appear as leaf nodes, as shown in Figure 2(b). Based on this graph representation, we employ the Graph Edit Distance (GED) (Abu-Aisheh et al., 2015), specifically the Tree Edit Distance (Zhang and Shasha, 1989), to measure the dissimilarity between a generated schema and the ground truth schema . GED represents the minimum cost required to transform into using a sequence of edit operations:
where is the set of all valid edit paths transforming to , and is the cost of an individual edit operation . Simply averaging raw GED scores across a dataset can be misleading. Smaller schemas naturally yield lower scores, possibly hiding significant relative errors, whereas larger schemas might skew the average. To mitigate this size bias, we normalize the GED by the maximum potential edit cost related to the ground truth schema’s size, defined as .
@@ 5.1 Models
To evaluate the capabilities of LLMs in shape generation, we compare against several baseline models. Our primary non-ML baseline is sheXer (Fernandez-Álvarez et al., 2022), a well-established system for KG shape extraction. For LLM-based comparisons, we selected gpt-4o-mini222https://platform.openai.com/docs/models/gpt-4o-mini and DeepSeek-V3 (DeepSeek-AI, 2024). Input data, including triples and global KG information, is retrieved using SPARQL queries against locally deployed SPARQL endpoints.
@@ 5.2 Prompt Engineering
To investigate how LLMs can effectively generate ShEx schemas and understand the impact of information type and generation mode, we designed experiments incorporating different types of information extracted from a KG. Based on the nature of this information, we developed and evaluated three distinct few-shot prompt engineering settings, summarized in Table 2.
@@
Inspired by findings that representative samples can suffice for shape extraction (Fernandez-Álvarez et al., 2022), this setting provides the LLM with a set of triples focused on the usage of specific predicates relevant to the entities. Different from the local setting, this setting focuses on a set of predicates and provides example triples where those predicates appear, potentially sampled across many different instances. The goal is to highlight common patterns for individual predicates. Similar to the local setting, the LLM generates the ShEx schema with few-shot examples.
| Models | Settings | YAGO Schema | Wikidata EntitySchema | ||||||||
| P | R | F1 | GED | NGED | P | R | F1 | GED | NGED | ||
| sheXer | / | 0.111 | 0.081 | 0.092 | 34.08 | 0.581 | 0.106 | 0.099 | 0.096 | 90.30 | 0.833 |
| gpt-4o-mini | local | 0.575 | 0.550 | 0.559 | 16.61 | 0.308 | 0.358 | 0.224 | 0.264 | 82.14 | 0.697 |
| global | 0.421 | 0.362 | 0.388 | 21.03 | 0.366 | 0.306 | 0.328 | 0.312 | 54.44 | 0.484 | |
| triples | 0.631 | 0.564 | 0.591 | 18.36 | 0.344 | 0.328 | 0.196 | 0.237 | 72.44 | 0.577 | |
| DeepSeek-V3 | local | 0.536 | 0.524 | 0.526 | 18.56 | 0.348 | 0.322 | 0.240 | 0.263 | 66.16 | 0.585 |
| global | 0.478 | 0.484 | 0.479 | 16.83 | 0.295 | 0.304 | 0.343 | 0.318 | 57.36 | 0.494 | |
| triples | 0.535 | 0.505 | 0.510 | 21.89 | 0.468 | 0.269 | 0.277 | 0.269 | 55.80 | 0.488 | |
| Settings | Matching Criteria | YAGO Schema | Wikidata EntitySchema | |||||
| Node Constraint | Cardinality | P | R | F1 | P | R | F1 | |
| local | Exact | Exact | 0.536 | 0.524 | 0.526 | 0.322 | 0.240 | 0.263 |
| Subclass | Exact | 0.551 | 0.540 | 0.542 | 0.413 | 0.303 | 0.335 | |
| Datatype | Exact | 0.634 | 0.625 | 0.626 | 0.537 | 0.396 | 0.437 | |
| Exact | Loosened | 0.589 | 0.582 | 0.580 | 0.366 | 0.280 | 0.303 | |
| Subclass | Loosened | 0.604 | 0.598 | 0.596 | 0.466 | 0.352 | 0.384 | |
| Datatype | Loosened | 0.753 | 0.751 | 0.746 | 0.613 | 0.464 | 0.507 | |
| global | Exact | Exact | 0.434 | 0.439 | 0.435 | 0.304 | 0.343 | 0.318 |
| Subclass | Exact | 0.521 | 0.528 | 0.523 | 0.482 | 0.550 | 0.507 | |
| Datatype | Exact | 0.698 | 0.709 | 0.701 | 0.577 | 0.655 | 0.606 | |
| Exact | Loosened | 0.441 | 0.446 | 0.442 | 0.394 | 0.451 | 0.415 | |
| Subclass | Loosened | 0.532 | 0.541 | 0.535 | 0.638 | 0.738 | 0.676 | |
| Datatype | Loosened | 0.751 | 0.764 | 0.755 | 0.793 | 0.910 | 0.839 | |
| triples | Exact | Exact | 0.535 | 0.505 | 0.510 | 0.269 | 0.277 | 0.269 |
| Subclass | Exact | 0.554 | 0.523 | 0.528 | 0.371 | 0.377 | 0.367 | |
| Datatype | Exact | 0.630 | 0.591 | 0.600 | 0.536 | 0.549 | 0.534 | |
| Exact | Loosened | 0.581 | 0.550 | 0.554 | 0.320 | 0.334 | 0.322 | |
| Subclass | Loosened | 0.604 | 0.570 | 0.575 | 0.449 | 0.461 | 0.448 | |
| Datatype | Loosened | 0.759 | 0.720 | 0.725 | 0.675 | 0.699 | 0.677 | |
@@ 5.3 Structured Generation
Even though LLMs excel at structured generation with well-defined JSON schemas, directly applying the full ShEx syntax to the LLM’s constraint decoding process is still challenging due to its complexity. To address this, our structured generation pipeline first simplifies the ShEx syntax within the context of the benchmark by converting it into a more manageable, JSON schema-like representation using ShExJ syntax. This process generates Pydantic models, each corresponding to a segment of ShEx syntax, to guide LLM constraint generation. We leverage the Instructor (Liu and Contributors, 2024) to enhance the LLM’s structured generation capabilities.
@@ 6 Results
Table 3 presents a comparative performance analysis of the baseline models on YAGOS and WES datasets. On YAGOS, gpt-4o-mini achieved the highest F1 (0.591) in its triples setting, while DeepSeek-V3’s global setting showed the best structural similarity (lowest NGED of 0.295). YAGOS’s generally stronger results are attributed to its ontology being largely derived and refined from schema.org, along with its smaller scale—having only half the average number of constraints and 80% of the average number of instances per class compared to WES. For the more challenging WES dataset with a complex predicate vocabulary and twice the constraints per class, DeepSeek-V3 in the global setting yielded the highest F1 (0.318). Both gpt-4o-mini and DeepSeek-V3 performed well on the NGED score in their global setting. Compared to the non-ML baseline sheXer, these results underscore the potent schema generation capabilities of LLM-based approaches.
@@
We further analyzed the performance of models under different matching criteria, using DeepSeek-V3 as an example (Table 4). Full results for other models are in Appendix C. In general, loosening the matching criteria leads to improved evaluation scores, as expected. Notably, when combining datatype abstraction with loosened cardinality, DeepSeek-V3 in the global setting achieved impressive F1 scores (0.755 on YAGO, 0.839 on WES). These high scores suggest that LLMs can generate schemas for certain practical validation scenarios, particularly where the primary goal is to ensure predicate completeness and general object constraints.
@@
As for individual criteria, allowing datatype abstraction provided the most significant performance gain, especially in the global setting. This indicates LLMs can effectively process given datatype information, finding it less challenging than inferring cardinality. Allowing approximate subclass matching improved classification scores more for WES compared with YAGOS, suggesting predicting the precise class list for referenced shapes is harder for WES, while generating related object classes is not. These findings indicate that although LLM-generated schemas may benefit from refinement, the required adjustments are generally less intensive or domain-specific than those for existing automated approaches, while still supporting effective validation.
@@
Figure 1 shows the distribution of results across five categories (correct and four error types) for models and settings on WES and YAGOS. Error types are: (1) missing predicates, (2) incorrect cardinality, (3) incorrect node constraint, and (4) both cardinality and node constraint are incorrect. Global settings effectively reduce missing predicates for models on both datasets. Missing predicates are more frequent on WES than YAGOS, likely due to WES’s larger set of candidate predicates. Global settings also tend to reduce instances of only incorrect node constraints, particularly on YAGOS. However, global settings often show a higher rate of errors where both cardinality and node constraint are incorrect, especially on the WES dataset.

### v2

@@ Abstract
Schemas play a vital role in ensuring data quality and supporting usability in the Semantic Web and natural language processing. Traditionally, their creation demands substantial involvement from knowledge engineers and domain experts. Leveraging the impressive capabilities of large language models (LLMs) in tasks like ontology engineering, we explore schema generation using LLMs. To bridge the resource gap, we introduce two datasets: YAGO Schema and Wikidata EntitySchema, along with novel evaluation metrics. The LLM-based pipelines utilize local and global information from knowledge graphs (KGs) to generate schemas in Shape Expressions (ShEx). Experiments demonstrate LLMs’ strong potential in producing high-quality ShEx schemas, paving the way for scalable, automated schema generation for large KGs. Furthermore, our benchmark introduces a new challenge for structured generation, pushing the limits of LLMs on syntactically rich formalisms.
@@ 1 Introduction
As crucial resources for quality assurance, KG schemas are generally classified into semantic, validating, and emergent schemas (Hogan et al., 2021). In this work, we focus on validating schemas, which are particularly effective in detecting structural and semantic inconsistencies and thereby ensuring data integrity (Gayo et al., 2018; Scherp et al., 2024). These schemas are defined through shapes, where each shape identifies a set of focus nodes within a graph and specifies the constraints and patterns they must adhere to (Ahmetaj et al., 2025). To represent such schemas, several dedicated languages have been developed, including W3C standards like Shapes Constraint Language (SHACL) (Knublauch and Kontokostas, 2017), Shape Expressions (ShEx) (Prud’hommeaux et al., 2014), and PG-Schema (Angles et al., 2023). However, developing high-quality validation schemas aligned with user needs remains a substantial challenge (Rabbani et al., 2022). Current automatic schema generation via pattern aggregation often inherits noise and errors from the source KGs (Rabbani et al., 2022). As a result, the primary approach involves manually writing schemas or refining those produced by basic pattern mining—a process that is both time-consuming and expensive, particularly for KGs with potentially millions of classes (Rabbani et al., 2022). The demonstrated success of LLMs in areas like ontology engineering (Lo et al., 2024; Zhang et al., 2025a) and graph-based reasoning (Wang et al., 2023) highlights their proficiency with structured data. Specifically, LLMs show aptitude for processing structured information, reasoning across graph and text modalities, and performing structured generation (He et al., 2024b; Jin et al., 2024). This aptitude strongly suggests their potential for automatically generating KG schemas.
Building on this, this paper investigates how LLMs can undertake this automatic schema generation for large KGs, especially where such schemas are missing or incomplete. Specifically, we address the research question: How can LLMs generate high-quality ShEx shapes for a given KG? We identify key challenges inherent to this task. First, due to the sheer scale of large KGs, pinpointing the essential information required for accurate ShEx generation by LLMs is difficult. In addition, even when relevant information is identified, investigating how LLMs can effectively leverage this structured data for schema generation is required. Moreover, the relative novelty and ongoing development of the ShEx standard means that publicly available ShEx corpora are scarce, potentially hindering LLM familiarity with the required syntax and conventions. To address these challenges, this paper makes the following main contributions:
@@
We introduce a benchmark comprising two datasets derived from large KGs—YAGO and Wikidata—along with associated metrics designed for comprehensive evaluation from diverse perspectives111The data is available at: https://doi.org/10.5281/zenodo.17128093.
We propose and evaluate pipelines designed to efficiently generate high-quality ShEx schemas by leveraging the structured generation capabilities of LLMs222The code is available at https://github.com/King-s-Knowledge-Graph-Lab/shapespresso.
@@ 3 Problem Formulation
Schema generation is a challenging task for both traditional rule-based methods (Rabbani et al., 2023) and ML approaches, due to several factors. Large KGs often suffer from quality issues, and pattern extraction may inadvertently propagate these issues into generated shapes. Moreover, the lack of benchmarks and high-quality ShEx ground truths makes training and evaluation of ML-based approaches particularly difficult.
@@ 4.1 Dataset
To address the need for benchmarks, we introduce two new dataset, YAGO Schema (YAGOS) and Wikidata EntitySchema (WES). Together, comprising 86 ShEx schema with a total of 2,552 constraints. Detailed statistics are presented in Table 1. Each schema in our datasets targets a specific class within the respective KG. While ShEx offers a rich syntax, the schemas in our datasets employ a community-prevalent subset to enhance simplicity and readability while remaining functional (detailed specification in Appendix B). In our setting, to simplify the syntax representation and ease evaluation and comparison, a shape schema includes a start shape that targets the focus class. The remaining shapes in the schema serve as references, each describing one or more classes that specify the range of objects allowed for certain predicates in the start shape. The schemas predominantly use four fundamental node constraint types: node kind constraints (e.g., IRI), datatype constraints (e.g., xsd:decimal, rdf:langString), value set constraints (often specifying object values, e.g., [schema:Organization]), and shape references (e.g., @<Person>, which links to another defined shape like <Person> { rdf:type [ schema:Person ] }).
The YAGOS dataset was constructed using YAGO 4.5 (Suchanek et al., 2024) as the input KG. Ground truth ShEx schema construction was informed by the existing SHACL schema and the official YAGO 4.5 design document. While the documentation provided a starting point, it covered only a subset of properties and constraints, so the final schemas were manually refined by retaining documented constraints and adding frequently used predicates identified through KG queries to match the intended scope.
The WES dataset was developed through a semi-automatic process. Target classes were selected from three sources: the community-curated Wikidata EntitySchema directory, classes mapped from YAGOS, and those mapped from Wikipedia categories. For each selected class, three knowledge engineers independently annotated its schema through an iterative process of automatic predicate compilation and prioritization, semantic refinement and deduplication, and cardinality and node constraint specification (datatype, value set, shape reference, or node kind). Constraints were retained if at least two annotators agreed, while disagreements were resolved through discussion or omission to ensure high-quality, consistent ShEx ground truth schemas. Further details of the dataset construction process are provided in Appendix A.
@@ 4.2.1 Similarity Metrics
Given that automatically generated schemas often require manual refinement (Rabbani et al., 2022), quantifying the structural similarity between a generated schema and its ground truth counterpart is crucial. To facilitate this comparison, we model each shape schema as a rooted, labeled tree graph. The shape label (or focus node) serves as the root, predicates form the first level of child nodes, node constraints linked to predicates occupy the next level, and cardinalities appear as leaf nodes, as shown in Figure 3(b). Based on this graph representation, we employ the Graph Edit Distance (GED) (Sanfeliu and Fu, 1983), specifically the Tree Edit Distance (Zhang and Shasha, 1989), to measure the dissimilarity between a generated schema and the ground truth schema . GED represents the minimum cost required to transform into using a sequence of edit operations:
where is the set of all valid edit paths transforming to , and is the cost of an individual edit operation . In this context, GED has a complexity of , where denotes the number of candidate constraints (i.e., distinct predicates), since schemas are represented as fixed-depth trees. Simply averaging raw GED scores across a dataset can be misleading. Smaller schemas naturally yield lower scores, possibly hiding significant relative errors, whereas larger schemas might skew the average. To mitigate this size bias, we normalize the GED by the maximum potential edit cost related to the ground truth schema’s size, defined as .
@@ 5.1 Models
To evaluate the capabilities of LLMs in shape generation and the effectiveness of different information types, we compare against several baseline models across multiple input settings. Our primary non-ML baseline is sheXer (Fernandez-Álvarez et al., 2022), a well-established system for KG shape extraction that relies on graph structure mining and directly takes RDF sources as input. We also include RDFShapeInduction (Mihindukulasooriya et al., 2018), a feature learning and machine learning (ML)-based approach, specifically to benchmark performance on cardinality prediction. For this baseline, we re-implemented the model using the provided feature lists and replaced its classification components with more advanced models, including random forests (Breiman, 2001) and gradient boosting (Friedman, 2001).
For LLM-based comparisons, we selected GPT-4o mini (Hurst et al., 2024) and DeepSeek-V3 (DeepSeek-AI, 2024) as backbone models. Figure 1 illustrates the workflow of our LLM-based generation pipeline. Given a target class, the pipeline retrieves the required information according to the chosen setting—including triples and global context—via SPARQL queries over the endpoints associated with the KG. To ensure scalability to large KGs, the approach first samples the subgraph of the target class to optimize information extraction. Across all settings, the input prompt lengths remain relatively stable regardless of KG size, and the number of retrieved triples is constrained to a manageable limit (up to 5).
@@ 5.2 Prompt Engineering
To investigate how LLMs can effectively generate ShEx schemas and understand the impact of information type and generation mode, we designed experiments incorporating different types of information extracted from a KG. Based on the nature of the information and existing KG-based prompt construction methods (Wen et al., 2024; Zhang et al., 2025b), we proposed three distinct few-shot prompt settings, summarized in Table 2.
@@
Inspired by findings that representative samples can suffice for shape extraction (Fernandez-Álvarez et al., 2022), this setting provides LLMs with a set of triples focused on the usage of specific predicates relevant to the entities. Different from the local setting, this setting focuses on a set of predicates and provides example triples where those predicates appear, potentially sampled across many different instances. The goal is to highlight common patterns for individual predicates. Similar to the local setting, the LLM generates the ShEx schema with few-shot examples.
@@ 5.3 Structured Generation
Structured generation enables LLMs to produce outputs adhering to precise formats, vital for applications like code generation (Ugare et al., 2024) and tool calling (Zhang et al., 2024). This involves generating token sequences that satisfy specified constraints, often defined by regular expressions (regex) or context-free grammars (CFGs). A key challenge is efficiently applying these constraints over the LLM’s large vocabulary without degrading speed or performance (Dong et al., 2024; Koo et al., 2024). Even though LLMs excel at structured generation with well-defined JSON schemas, directly applying the full ShEx syntax to the LLM’s constraint decoding process is still challenging due to its complexity. To address this, our structured generation pipeline first simplifies the ShEx syntax within the context of the benchmark by converting it into a more manageable, JSON schema-like representation using ShExJ syntax. This process generates Pydantic models, each corresponding to a segment of ShEx syntax, to guide LLM constraint generation. We leverage the Instructor (Liu and Contributors, 2024) to enhance the LLM’s structured generation capabilities.
@@ 6 Results
Models Settings YAGOS WES P R F1 GED NGED P R F1 GED NGED sheXer / 0.111 0.081 0.092 34.08 0.581 0.106 0.099 0.096 90.30 0.833 GPT-4o mini local 0.575 0.550 0.559 16.61 0.308 0.358 0.224 0.264 82.14 0.697 global 0.421 0.362 0.388 21.03 0.366 0.306 0.328 0.312 54.44 0.484 triples 0.631 0.564 0.591 18.36 0.344 0.328 0.196 0.237 72.44 0.577 DeepSeek-V3 local 0.536 0.524 0.526 18.56 0.348 0.322 0.240 0.263 66.16 0.585 global 0.478 0.484 0.479 16.83 0.295 0.304 0.343 0.318 57.36 0.494 triples 0.535 0.505 0.510 21.89 0.468 0.269 0.277 0.269 55.80 0.488
Settings Matching Criteria WES Node Constraint Cardinality P R F1 local Exact Exact 0.322 0.240 0.263 Subclass Exact 0.413 0.303 0.335 Datatype Exact 0.537 0.396 0.437 Exact Loosened 0.366 0.280 0.303 Subclass Loosened 0.466 0.352 0.384 Datatype Loosened 0.613 0.464 0.507 global Exact Exact 0.304 0.343 0.318 Subclass Exact 0.482 0.550 0.507 Datatype Exact 0.577 0.655 0.606 Exact Loosened 0.394 0.451 0.415 Subclass Loosened 0.638 0.738 0.676 Datatype Loosened 0.793 0.910 0.839 triples Exact Exact 0.269 0.277 0.269 Subclass Exact 0.371 0.377 0.367 Datatype Exact 0.536 0.549 0.534 Exact Loosened 0.320 0.334 0.322 Subclass Loosened 0.449 0.461 0.448 Datatype Loosened 0.675 0.699 0.677
Model WES Acc (min) Acc (max) Acc sheXer 0.982 0.402 0.394 GPT-4o mini (local) 0.988 0.441 0.432 GPT-4o mini (global) 0.900 0.584 0.520 GPT-4o mini (triples) 0.988 0.467 0.457 DeepSeek-V3 (local) 0.975 0.490 0.474 DeepSeek-V3 (global) 0.978 0.499 0.485 DeepSeek-V3 (triples) 0.933 0.547 0.517 RDFShapeInduction (DT) 0.992 0.566 0.559 RDFShapeInduction (MLP) 0.988 0.606 0.596 RDFShapeInduction (RF) 0.994 0.645 0.640 RDFShapeInduction (GB) 0.993 0.662 0.656
Model WES P R F1 GED NGED sheXer 0.083 0.071 0.069 116.40 0.849 GPT-4o mini (local) 0.329 0.139 0.188 104.60 0.704 GPT-4o mini (global) 0.239 0.290 0.259 80.50 0.609 GPT-4o mini (dt-global) 0.273 0.259 0.264 66.50 0.474 GPT-4o mini (gb-global) 0.283 0.327 0.301 67.30 0.500 GPT-4o mini (triples) 0.301 0.170 0.205 96.10 0.611 DeepSeek-V3 (local) 0.285 0.192 0.221 86.10 0.586 DeepSeek-V3 (global) 0.232 0.290 0.255 84.70 0.637 DeepSeek-V3 (dt-global) 0.284 0.272 0.275 61.70 0.441 DeepSeek-V3 (gb-global) 0.302 0.350 0.321 60.60 0.456 DeepSeek-V3 (triples) 0.254 0.250 0.244 81.50 0.601
@@
Table 3 presents a comparative performance analysis of the baseline models on YAGOS and WES datasets. On YAGOS, GPT-4o mini achieved the highest F1 (0.591) in its triples setting, while DeepSeek-V3’s global setting showed the best structural similarity (lowest NGED of 0.295). YAGOS’s generally stronger results are attributed to its ontology being largely derived and refined from schema.org, along with its smaller scale—having only half the average number of constraints and 80% of the average number of instances per class compared to WES. For the more challenging WES dataset with a complex predicate vocabulary and twice the constraints per class, DeepSeek-V3 in the global setting yielded the highest F1 (0.318). Both GPT-4o mini and DeepSeek-V3 performed well on the NGED score in their global setting. Compared to the non-ML baseline sheXer, these results underscore the potent schema generation capabilities of LLM-based approaches.
@@
We further analyze the performance of models under different matching criteria, using DeepSeek-V3 on the WES dataset as an example (Table 4). Full results for other models are in Appendix C. In general, loosening the matching criteria leads to improved evaluation scores, as expected. Notably, when combining datatype abstraction with loosened cardinality, DeepSeek-V3 in the global setting achieved impressive F1 scores (0.839). These results suggest that LLMs can generate schemas for certain practical validation scenarios, particularly where the primary goal is to ensure predicate completeness and general object constraints.
@@
As for individual criteria, allowing datatype abstraction provides the most significant performance gain, especially in the global setting. This indicates LLMs can effectively process given datatype information, finding it less challenging than inferring cardinality. Allowing approximate subclass matching improved classification scores more for WES compared with YAGOS, suggesting predicting the precise class list for referenced shapes is harder for WES, while generating related object classes is not. These findings indicate that although LLM-generated schemas may benefit from refinement, the required adjustments are generally less intensive or domain-specific than those for existing automated approaches, while still supporting effective validation.
@@
Figure 2 shows the distribution of results across five categories (correct and four error types) for models and settings on WES and YAGOS. Error types are: (1) missing predicates, (2) incorrect cardinality, (3) incorrect node constraint, and (4) both cardinality and node constraint are incorrect. Global settings effectively reduce missing predicates for models on both datasets. Missing predicates are more frequent on WES than YAGOS, likely due to WES’s larger set of candidate predicates. However, global settings show a higher rate of cases where both cardinality and node constraint are incorrect, especially on the WES dataset.
Specifically for cardinality prediction, we compare LLM results with RDFShapeInduction using different ML models based on the accuracy of the lower limit, upper limit, and their combination (Table 5). ML models (e.g., gradient boosting) surpass LLMs, indicating that LLMs still have room for improvement on this subtask. To leverage this, we replace cardinality prediction components with ML models in LLM-based generation pipelines under the global setting. The results on the WES dataset are shown in Table 6, where LLMs augmented with ML models outperform LLMs alone. Similar improvements are observed on the YAGOS dataset (Tables 9 and 10, Appendix C). The strong performance of traditional ML models suggests that incorporating features learned from these models into LLM inputs is a promising strategy for improving end-to-end schema generation.